OpenStudy (anonymous):

what is 10ft 8in - 3ft 11in

OpenStudy (anonymous):

I think we have to go like: 12*10=120+8=128 12*3=36+11=47. 128-47=81. I think this is the answer.

OpenStudy (anonymous):

6 feet 9 inches :)